Grade 11Physical Chemistry

Calculate the volume of o2 required for complete combustion of 1lit ch4 gas at NTP.

Ankit kumar
8 Years agoGrade 11

1 Answer

shonit
8 Years ago
CH4[g] + 2 O2[g] -> CO2[g] + 2 H2O[g] + energy for one mole of methane 2 mole of o2 required, so for 1/22.7=0.044 of methane , 0.088moles of o2 required ,volume=0.088*22.7=1.99276